OpenStudy (anonymous):

Rewrite y =(x2 + 14x) + 29 in general form. (1 point)

OpenStudy (ybarrap):

how do you interpret "general form?"

OpenStudy (anonymous):

taking it from standard to general form

OpenStudy (ybarrap):

what is that form, symbolically?

OpenStudy (anonymous):

y=a(x-h)^2+k

OpenStudy (ybarrap):

Using http://www.mathsisfun.com/algebra/line-equation-general-form.html, general form is Ax + By + C = 0

OpenStudy (anonymous):

oh ok

OpenStudy (ybarrap):

HTH

OpenStudy (anonymous):

by completing to square?

OpenStudy (anonymous):

the*

OpenStudy (ybarrap):

General forms are for straight lines, this is not a straight line if it is a quadratic

OpenStudy (ybarrap):

For quadratics, General form is Ax^2+Bx+C=0

OpenStudy (ybarrap):

For general form of quadratics - http://en.wikipedia.org/wiki/Quadratic_equation

OpenStudy (anonymous):

y=(x^2+14x+__)+29-__

OpenStudy (ybarrap):

x^2 + 14x + 29 - This is in General Form

OpenStudy (ybarrap):

brb

OpenStudy (ybarrap):

Does this make sense, General Form is not in any factoring form and separated by degrees with constant coefficients

OpenStudy (anonymous):

this what my teacher said, General form is y=a(x-h)^2+k. We use the process of completing the square to take it from standard to general form. Group the first two terms and add in blanks for where you will complete the square y=(x^2+14x+__)+29-__ Whatever we add we must also subtract so that the equation stays balanced.

OpenStudy (abb0t):

was this taken from an exam? it says "(1 point)" and posting exam questions is not allowed. Read CoC.

OpenStudy (anonymous):

m=no its a practice assignment that im resubmiting

OpenStudy (ybarrap):

Then you should complete the square and that technique will allow you to get to that form, whatever they want to call it.